The flower garden has the shape of a right traingle. 51 ft. of a perennial border forms the hypotenuse of the traingle, and one leg is 21 ft. longer than the other leg. Find the lengths of the legs. What are the lengths of the legs in ft?
Answer by checkley77(12844)

X^2+(X+21)^2=51^2
X^2+X^2+42X+441=2,601
2X^2+42X+441-2,601=0
2X^2+42X-2,160=0
2(X^2+21X-1,080)=0
2(X-24)(X+45)=0
X-24=0
X=24 ANS.
X+45=0
X=-45 ANS.